OpenStudy (anonymous):

In the system: x + y = 10 and x - 2y = 8, Marie said that to eliminate x the first step would be to multiply the first equation by 2. Is she correct? If not, why not?

OpenStudy (anonymous):

cause we have to eliminate x and it can b eliminated without multiplying by 2... just subtract 2nd equation from the first one..

OpenStudy (anonymous):

\[=>(x+y=10)-(x-2y=8) \]

OpenStudy (anonymous):

\[=> 4y=2\]

OpenStudy (anonymous):

so is she correct

OpenStudy (anonymous):

sorry 3y=2

OpenStudy (anonymous):

y=2/3...

OpenStudy (anonymous):

wait shes wrong because you dont have to multiply right

OpenStudy (anonymous):

now check did we multiply by 2 in frst step or not??? for eliminating x??

OpenStudy (anonymous):

Shes right.

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

nope.. she z wrng

OpenStudy (anonymous):

oh okay thanks

OpenStudy (anonymous):

v jxt eliminate by subtraction.. rite??

OpenStudy (anonymous):

(2x+2y=20) + (x -2y=8) =3x=28

OpenStudy (anonymous):

you add since the second 2y is negative.